Okay how did I find the area of an enclosed rectangle if the entire fence length is 820ft and one side 118ft?
Answer by Theo(13342)

P = 2L + 2W

L = length
W = width
P = perimeter

L = 118
P = 820

P = 2L + 2W becomes:

820 = 2*118 + 2W

solve for W to get W = (820 - 236) / 2 = 292

L = 118
W = 292

A = area

A = L * W = 118 * 292 = 34456 square feet.